Friday, 21 August

Plenary Lectures

9:30 – 10:30  Auditorium Maximum H 105

Helmut H.W. Hofer, Courant Institute, New York University, New York, USA Dynamics, Topology, and Holomorphic Curves (I: p.8)

Chair: Vladimir Arnold, Steklov Mathematical Institute, Moscow, Russia

11:00 – 12:00  Auditorium Maximum H 105

Vladimir Voevodsky, Northwestern University, Evanston, USA Homotopy Theory of Algebraic Varieties (I: p.28)

Chair: Eric Mark Friedlander, Northwestern University, Evanston, USA

ICM’98 – Program  1